A NEW ATLAS OF THE DIFFUSE INTERSTELLAR BANDS: HD 183143

LEWIS M. HOBBS, University of Chicago, Yerkes Observatory, Williams Bay, WI 53191; DON-ALD G. YORK, University of Chicago, Chicago, IL 60637; JULIE A. THORBURN, Carthage College, Kenosha, WI 53140; THEODORE P. SNOW, University of Colorado, Boulder, CO 80309; MICHAEL BISHOF, University of Chicago, Chicago, IL 60637; SCOTT D. FRIEDMAN, Space Telescope Science Institute, Baltimore, MD 21218; \underline{BENJAMIN J. McCALL}, University of Illinois at Urbana-Champaign, Urbana, IL 61801; TAKESHI OKA, University of Chicago, Chicago, IL 60637; BRIAN L. RACHFORD, Embry-Riddle Aeronautical University, Prescott, AZ 86301; PAULE SON-NENTRUCKER, Johns Hopkins University, Baltimore, MD 21218; DANIEL E. WELTY, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

We present our second catalog^{a} of the diffuse interstellar bands (DIBs), based on high signal-to-noise ratio echelle spectra of HD 183143 obtained at the Apache Point Observatory. This catalog complements our first catalog,^{b} which was based on spectra of HD 204827. Unlike the sightline towards HD 204827, which hosts a high column density of C_{2} and C_{3}, the sightline towards HD 183143 has no detectable amount of carbon chain molecules and therefore samples a somewhat chemically distinct environment.

Our catalog of HD 183143 contains 414 DIBs, of which 135 (or 33%) were not reported in four previous modern DIB surveys. When combined with our catalog of HD 204827, the total number of distinct DIBs observed is ∼545. Our collaboration’s website at http://dibdata.org contains a complete listing of the properties of the observed DIBs in tabular format, PDF files containing the observed spectra, and an interactive spectral plotting tool that enables users to rescale the spectra. Our hope is that this dataset will facilitate the comparison of laboratory molecular spectra with the astronomical observations, and ultimately the identification of the molecular carriers of the DIBs.

aL. M. Hobbs et al., Astrophysical Journal 705, 32-45 (2009)

bL. M. Hobbs et al., Astrophysical Journal 680, 1256-1270 (2008)